# DeFoG: Discrete Flow Matching for Graph Generation

### v1

@@ 6.2 Molecular Graph Generation
Table 2: Large molecule generation performance. Only iterative denoising-based methods are reported here. Respecive full versions in Tab. 9 (MOSES) and Tab. 10 (Guacamol), Sec. F.2.
MOSES Guacamol
[…]
Training set 100.0 100.0 0.0 100.0 0.01 0.64 99.1 100.0 100.0 0.0 99.9 92.8
DiGress (Vignac et al., 2022) 85.7 100.0 95.0 97.1 1.19 0.52 14.8 85.2 85.2 85.1 92.9 68.0
DisCo (Xu et al., 2024) 88.3 100.0 97.7 95.6 1.44 0.50 15.1 86.6 86.6 86.5 92.6 59.7
Cometh (Siraudin et al., 2024) 90.5 99.9 92.6 99.1 1.27 0.54 16.0 98.9 98.9 97.6 96.7 72.7
DeFoG (# steps = 50) 83.9 99.9 96.9 96.5 1.87 0.50 23.5 91.7 91.7 91.2 92.3 57.9
DeFoG (# steps = 500) 92.8 99.9 92.1 98.9 1.95 0.55 14.4 99.0 99.0 97.9 97.7 73.8
@@
Molecular design is a prominent real-world application of graph generation. We evaluate DeFoG’s performance on this task using the QM9 (Wu et al., 2018), MOSES (Polykovskiy et al., 2020), and Guacamol (Brown et al., 2019) datasets. For QM9, we follow the dataset split and evaluation metrics from Vignac et al. (2022), presenting the results in Sec. E.1.2, Tab. 8. For the larger MOSES and Guacamol datasets, we adhere to the training setup and evaluation metrics established by Polykovskiy et al. (2020) and Brown et al. (2019), respectively, with results in Tabs. 9 and 10.
As illustrated in Tab. 2, DeFoG outperforms existing diffusion models. It achieving state-of the Validity while preserving high uniqueness on MOSES. On Guacamol, it consistently ranks best, followed by Cometh, another work utilizing a continuous-time framework. Notably, DeFoG approaches the performance of existing diffusion models with only 10% of the sampling steps. This result is further investigated in the following section.
@@ 6.3 Efficiency Improvement
As highlighted in Tabs. 1, 2 and 8, DeFoG attains similar performance, and in some cases even outperforms, several diffusion-based methods with just 5% or 10% of their sampling steps. This efficiency is a result of DeFoG’s sampling stage flexibility, whose optimization is enabled due to the flexible and disentangled training-sampling procedure within DFM. To further demonstrate the advantages of sampling optimization, Figure 3(a) shows the cumulative effect of each optimization step discussed in Sec. 4.2.2. As illustrated, starting from a vanilla DeFoG model, which initially performs slightly worse than DiGress, we sequentially incorporate algorithmic improvements, with each stage indicated by an increasing number of $+$ symbols. Each addition leads to a progressive improvement over the previous stage across both the Planar and QM9 datasets under various numbers of steps for sampling, with the final performance significantly surpassing the vanilla model using only 50 steps on the Planar dataset. While the benefits of each optimization component may vary across datasets, the sampling optimization pipeline remains highly efficient, enabling quick hyperparameter exploration without requiring retraining, which is a significantly more resource-intensive stage (see Sec. E.2). We provide more details on the impact of each sampling optimization strategy across datasets in Sec. A.4.
@@
While our primary focus is on optimizing the sampling process given its computational efficiency, we emphasize that DeFoG’s training optimization strategies can further push its performance. To illustrate this, Figure 3(b) shows the convergence curves for the tree and MOSES datasets. We first observe that, with the same model, leveraging only sampling distortion can further enhance its performance beyond the vanilla implementation. This suggests that optimizing the sampling procedure can be particularly useful in settings where computational resources are limited and the model is undertrained, as detailed in Sec. A.6. Moreover, when an appropriate sample distortion is known (e.g., polydec distortion is shown to be particularly preferred across molecular datasets), applying it to both train distortion and sample distortion typically further improves performance. Although DeFoG achieves faster convergence with both distortions, we remark that DiGress also shows good convergence speed on the Tree dataset, due to the well-tuned noise schedule employed for joint training and sampling. The mutual impact of training and sampling distortion is discussed in Sec. B.2. Besides, in the same section, we discuss a heuristic on which distortion to use for each dataset according to DeFoG’s vanilla dynamics. Further details on how the initial distribution affects training efficiency are provided in Sec. B.1.
@@
Figure 3: DeFoG’s improvements on sampling and training efficiency.
6.4 Conditional Generation
Setup
Tertiary Lymphoid Structure (TLS) graph datasets have been recently released with graphs built from digital pathology data (Madeira et al., 2024). These graphs are split into two subsets — low TLS and high TLS — based on their TLS content, a biologically informed metric reflecting the cell organization in the graph structure. Previously, models were trained and evaluated separately for each subset. To demonstrate the flexibility of DeFoG, we conditionally train it across both datasets simultaneously, using the low/high TLS content as a binary label for each graph. More details on the used conditional framework in Appendix D.
Table 3: TLS conditional generation results.
TLS Dataset
Model V.U.N. $\uparrow$ TLS Val. $\uparrow$
Train set 0.0 100
GraphGen (Goyal et al., 2020) 40.2±3.8 25.1±1.2
BiGG (Dai et al., 2020) 0.6±0.4 16.7±1.6
SPECTRE (Martinkus et al., 2022) 7.9±1.3 25.3±0.8
DiGress+ (Madeira et al., 2024) 13.2±3.4 12.6±3.0
ConStruct (Madeira et al., 2024) 99.1±1.1 92.1±1.3
DeFoG (# steps = 50) 44.5±4.2 93.0±5.6
DeFoG (# steps = 1,000) 94.5±1.8 95.8±1.5
Results
@@ 7 Conclusion
We introduce DeFoG, a flexible discrete flow-matching framework for graph generation. Our theoretical contributions ensure that the denoising process preserves the graph distribution. Extensive experiments show that DeFoG achieves state-of-the-art results across tasks such as synthetic, molecular, and TLS graph generation, demonstrating real-world applicability and efficiency with just $50$ sampling steps. Although limited computational resources restricted a thorough hyperparameter search, leaving room for further optimization, DeFoG’s performance highlights its potential. While our work focuses on graphs, the flexibility of DeFoG and the proposed theoretical guarantees suggest that some of our techniques may generalize to other data modalities, opening interesting avenues for future research.
Reproducibility Statement
Code is available at github.com/manuelmlmadeira/DeFoG. We include detailed descriptions of the datasets, experimental setup and methodologies used in Secs. 6 and E. In particular, the specific hyperparameters employed in our experiments are documented in this paper in Sec. E.3 and in the code repository. The full version and proofs of theoretical results can be found in Appendix C.
@@
Acknowledgements
Appendix
[…]
Appendix A Sample Optimization
@@
In this section, we explore the proposed sampling optimization in more detail. We start by analysing the different time distortion functions in Sec. A.1. Next, in Sec. A.2, we prove that the proposed target guidance mechanism actually satisfies the Kolmogorov equation, thus yielding valid rate matrices and, in Sec. A.3, we provide more details about the detailed balance equation and how it widens the design space of rate matrices. In Sec. A.4, we also describe the adopted sampling optimization pipeline. Finally, in Sec. A.5, we provide more details to better clarify the dynamics of the sampling process.
A.1 Time Distortion Functions
@@
Note that even though we apply the same time distortions for both training and sample stages, in each setting they have different objectives: in training, the time distortions skew the PDFs from where $t^{\prime}$ is sampled, while in sampling they vary the step sizes.
A.2 Target Guidance
@@
A.3 Detailed Balance, Prior Incorporation, and Stochasticity
Campbell et al. (2024) show that although their $z_{1}$-conditional formulation of $R^{*}t$ generates $p{t|1}$, it does not span the full space of valid rate matrices — those that satisfy the conditional Kolmogorov equation (Eq. 11). They derive sufficient conditions for identifying other valid rate matrices. Notably, they demonstrate that matrices of the form

### v3

@@ 4 Related Work
Table 2: Large molecule generation performance. Only iterative denoising-based methods are reported here. Respective full versions in Tab. 10 (Guacamol) and Tab. 9 (MOSES), Sec. G.3.
Guacamol MOSES
Model Val. $\uparrow$ V.U. $\uparrow$ V.U.N.$\uparrow$ KL div $\uparrow$ FCD $\uparrow$ Val. $\uparrow$ Unique. $\uparrow$ Novelty $\uparrow$ Filters $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Scaf $\uparrow$
Training set 100.0 100.0 0.0 99.9 92.8 100.0 100.0 0.0 100.0 0.01 0.64 99.1
DiGress (Vignac et al., 2022) 85.2 85.2 85.1 92.9 68.0 85.7 100.0 95.0 97.1 1.19 0.52 14.8
DisCo (Xu et al., 2024) 86.6 86.6 86.5 92.6 59.7 88.3 100.0 97.7 95.6 1.44 0.50 15.1
Cometh (Siraudin et al., 2024) 98.9 98.9 97.6 96.7 72.7 90.5 99.9 92.6 99.1 1.27 0.54 16.0
DeFoG (10% steps) 91.7 91.7 91.2 92.3 57.9 83.9 99.9 96.9 96.5 1.87 0.50 23.5
DeFoG 99.0 99.0 97.9 97.7 73.8 92.8 99.9 92.1 98.9 1.95 0.55 14.4
5 Experiments
This section highlights DeFoG’s SOTA performance, enabled by its highly decoupled framework and effective sampling methods. We present DeFoG’s performance in generating graphs with diverse topological structures and in molecular datasets with rich prior information. We also provide ablations to showcase the effectiveness and necessity of each proposed sampling method. We highlight the best result and the second-best method. Results for generation with 5-10% of steps used by previous SOTA diffusion models are also provided to demonstrate DeFoG’s sampling efficiency. We show DeFoG’s versatility on conditional generation for digital pathology in Sec. G.1.  [^1] Code at github.com/manuelmlmadeira/DeFoG.
Setup
To isolate the effect of the network architecture, we build DeFoG on the best-performing graph transformer on generative tasks (Vignac et al., 2022). To enhance expressivity, we incorporate Relative Random Walk Probabilities (RRWP) (Ma et al., 2023; Siraudin et al., 2024) as node and edge features. More details on the architecture in Sec. F.1. Our ablations in Sec. G.4 show that, while RRWP features encode structural properties more efficiently and effectively than prior alternatives, DeFoG’s disentangled framework is the primary driver of performance gains. Importantly, the overall architecture, together with RRWP features, is guaranteed to be permutation equivariant (see Sec. D.2.1).
5.1 Graph Generation Performance
Synthetic Graph Generation
We evaluate DeFoG using the widely adopted Planar, SBM (Martinkus et al., 2022), and Tree datasets (Bergmeister et al., 2023), along with the associated evaluation methodology. In Tab. 1, we report the proportion of generated graphs that are valid, unique, and novel (V.U.N.), as well as the average ratio of the usual distances between graph statistics of the generated and test sets relative to the train and test sets (Ratio) to assess sample quality. As shown in Tab. 1, for the Planar dataset, DeFoG achieves the best performance across both metrics, with a nearly saturated V.U.N. value of 99.5%. On the Tree dataset, it is only surpassed by HSpectre, which leverages a local expansion procedure particularly well-suited to hierarchical structures like trees. On the SBM dataset, DeFoG achieves the highest V.U.N. score among all methods, even with just 50 steps.
Molecular Graph Generation
@@ 5.1 Graph Generation Performance
Molecular design is a prominent real-world application of graph generation. We evaluate DeFoG’s performance on this task using the QM9 (Wu et al., 2018), ZINC250k (Sterling & Irwin, 2015), MOSES (Polykovskiy et al., 2020), and Guacamol (Brown et al., 2019) datasets. For QM9, we follow the dataset split and evaluation metrics from Vignac et al. (2022), presenting the results in Sec. F.2.2, Tab. 8. For ZINC250k, we provide the experimental setup in Appendix F. For the larger MOSES and Guacamol datasets, we adhere to the training setup and evaluation metrics established by Polykovskiy et al. (2020) and Brown et al. (2019), respectively, with results in Tabs. 9 and 10. As illustrated in Tab. 2, on Guacamol, it ranks best across all metrics, achieving a nearly saturated validity of 99.0%. Notably, DeFoG achieves over 90% validity with just 10% of the sampling steps, surpassing the well-established baseline DiGress with 500 steps. On MOSES, DeFoG also outperforms diffusion models, achieving SOTA validity of 92.8% while maintaining a high uniqueness of 99.9%.
@@ 5.2 Efficiency Improvement
We now show that DeFoG enhances both training and sampling efficiency significantly across diverse datasets.
@@
Figure 4(a) highlights the cumulative benefits of sampling approaches from Sec. 3.2. Starting with a vanilla DeFoG model (initially slightly below DiGress), each optimization step, denoted by $+$ symbols, progressively improves performance, culminating in significant gains using only 50 steps on the Planar dataset. This demonstrates that the three sampling approaches are each essential components for optimizing generation.
@@
Figure 4(b) illustrates convergence curves for the tree and MOSES datasets. We observe that incorporating sampling distortion enhances performance significantly beyond the vanilla implementation, making it particularly useful for generation with undertrained models in resource-constrained settings (see Sec. B.6). Additionally, applying the same optimal distortion found in sampling during training typically yields further gains in convergence (see Sec. C.2). The convergence on some graph datasets may also benefit from an appropriate initial distribution, as shown for SBM in Sec. C.1.
@@
Figure 4: DeFoG’s improvements on efficiency.
5.3 Ablations
(a) Sample Distortion.
(b) Target guidance.
(c) Stochasticity.
Figure 5: Effect of proposed sampling approaches on the Planar dataset. Higher values on the vertical axis correspond to more favorable values for both V.U.N. and Ratio.
Here, we focus on evaluating the impact of different sampling approaches. We start from the vanilla sampling setup and sweep over sample distortion, target guidance, or stochasticity independently. More details in Sec. B.4. For illustration, here we focus on the Planar dataset:
Sample Distortion
In Figure 5(a), we observe that cos and polydec distortions which emphasize refinement at later steps, perform better on the Planar dataset. This aligns with the intuition that, unlike continuous data undergoing gradual refinement, graphs often experience abrupt transitions due to the random sampling of categorical data. These transitions can violate hard constraints, such as planarity, as categorical values shift abruptly (e.g., from 00 to $1$ in one-hot encoding) when $t$ approaches 1. These later steps are thus critical for error detection and correction. On the contrary, for datasets like SBM, where properties are not deterministic and such strict constraints are absent, this refinement does not provide any advantage (see Sec. B.1).
Target Guidance
As shown in Figure 5(b), $\omega$ improves both V.U.N. and Ratio by biasing generation toward predicted clean data. However, excessive $\omega$ skews the generated distribution to the high density regions of training set distribution, leading to higher planarity (reflected by V.U.N.) but increased divergence from test graphs (reflected by Ratio). We also observe that the Ratio only begins to worsen at $0.1$ with 50 steps, compared to $0.02$ with 1000 steps. This demonstrates that target guidance is particularly effective with fewer steps, where the generation process becomes more challenging due to larger transitions, as it steers the model toward higher-confidence regions, safeguarding generative performance.
Stochasticity
As shown in Figure 5(c), a moderate level of stochasticity benefits both metrics, while extreme values introduce excessive noise, disrupting the generation process. This indicates a sweet spot exists between effective error correction and over-stochasticity. Furthermore, the V.U.N. of generated graphs decreases with increasing $\eta$ values when more steps are utilized (drop after $\eta=100$ for 1000 steps vs. $\eta=10$ for 50 steps). As this approach preserves the Kolmogorov equation, it benefits from more sampling steps to mitigate simulation errors.
@@ 6 Conclusion
We introduce DeFoG, a novel discrete flow matching framework for graphs. This formulation enables training-sampling decoupling, which we ground theoretically to ensure faithful graph distribution modeling. Extensive experiments demonstrate the importance of our proposed strategies in achieving state-of-the-art performance on synthetic and molecular graph generation tasks. DeFoG currently employs a simple but efficient hyperparameter search for sampling, yielding impressive results and underscoring its potential for further improvement with more advanced search algorithms. Future work will further explore purely sampling-stage methods to enhance performance. Generating high-quality graphs in even fewer steps and scaling to larger graphs also remain key challenges.
Acknowledgements
@@
Appendix Overview
In the Appendix, we provide additional details organized as follows:
  1. 1.
    Appendix A: Contextualizing Related Research.
  2. 2.
    Appendix B: Sample Optimization.
  3. 3.
    Appendix C: Train Optimization.
  4. 4.
    Appendix D: Theoretical Results.
  5. 5.
    Appendix E: Conditional Generation.
  6. 6.
    Appendix F: Experimental Details.
  7. 7.
    Appendix G: Additional Results.
Appendix A Contextualizing Related Research
In this section, we further contextualize DeFoG within the scope of related work. We begin by introducing the methods used for comparison with DeFoG in Sec. A.1. Subsequently, we outline the key distinctions between DeFoG and existing diffusion-based graph generative models in Sec. A.2.
A.1 Overview of Compared Methods
In Sec. 5, we evaluate DeFoG against a diverse set of graph generative models, which we introduce below:
  • GraphRNN (You et al., 2018) and GRAN (Liao et al., 2019), two pioneering autoregressive models for graph generation;
  • SPECTRE (Martinkus et al., 2022), a spectrally conditioned GAN-based model for graph generation;
  • DiGress (Vignac et al., 2022), the first discrete diffusion model for graph generation;
  • EDGE (Chen et al., 2023), a discrete diffusion model leveraging graph sparsity and degree guidance for scalability.
  • BwR (Diamant et al., 2023), which focuses on efficient graph representations via bandwidth restriction schemes that are compatible with various graph generation models. We report its results in combination with EDP-GNN (Niu et al., 2020), which was the first graph diffusion model;
  • BiGG (Dai et al., 2020), an autoregressive model that exploits graph sparsity and training parallelization to scale to larger graphs;
  • GraphGen (Goyal et al., 2020), a scalable autoregressive approach utilizing graph canonization with minimum DFS codes, notable for being domain-agnostic and inherently supporting attributed graphs;
  • HSpectre (Bergmeister et al., 2023), a hierarchical graph generation method that utilizes a score-based formulation for iterative local expansion steps;
  • DisCo (Xu et al., 2024) and Cometh (Siraudin et al., 2024), two continuous-time discrete diffusion models for graph generation;
  • GruM (Jo et al., 2024), which employs a diffusion mixture to explicitly learn the final graph topology and structure;
  • CatFlow (Eijkelboom et al., 2024), which results from the instantiation of variational flow matching to graph generation.
A.2 DeFoG and Graph Diffusion Models
In this section, we contextualize DeFoG in relation to existing graph diffusion models.
A.2.1 From Continuous to Discrete State-spaces
Early diffusion-based graph generative models extended continuous diffusion and score-based methods from image generation to graphs by relaxing adjacency matrices into continuous state-spaces (Niu et al., 2020; Jo et al., 2022). However, this approach overlooks the inherent discreteness of graph-structured data, resulting in topologically uninformed noising processes. For instance, these methods often destroy graph sparsity and generate noisy complete graphs (Vignac et al., 2022; Xu et al., 2024; Siraudin et al., 2024), making it more challenging for denoising neural networks to recover meaningful structural properties from the noisy inputs. Some recent formulations operating on continuous state-spaces have tried to overcome these limitations: GruM (Jo et al., 2024) introduces an endpoint-conditioned diffusion mixture strategy to enhance accuracy by explicitly learning final graph structures, while CatFlow (Eijkelboom et al., 2024) proposes variational flow matching to handle categorical data more effectively.
Alternatively, discrete diffusion models have emerged as a more natural solution, directly preserving the discrete nature of graph data (Vignac et al., 2022; Haefeli et al., 2022). These models have demonstrated state-of-the-art performance across a variety of applications, including neural architecture search (Asthana et al., 2024), combinatorial optimization (Sun & Yang, 2024), molecular generation (Irwin et al., 2024), and reaction pathway design (Igashov et al., 2024).
DeFoG aligns with this second family of methods, modeling nodes and edges in discrete state-spaces to leverage the structural properties of graph data effectively.
A.2.2 From Discrete to Continuous time
The initial discrete-time diffusion frameworks for graph generation (Vignac et al., 2022; Haefeli et al., 2022) were built upon Discrete Denoising Diffusion Probabilistic Models (D3PMs) (Austin et al., 2021), which operate with a fixed partitioning of time. This discretization constrains the model to denoise at specific time points and ties the sampling process to the same fixed time steps used during training, leading to a rigid coupling between the training and sampling stages. Such inflexibility in time discretization can limit the quality of generated graphs.
In contrast, continuous-time discrete diffusion frameworks (Campbell et al., 2022; Sun et al., 2023) overcome these limitations by enabling the model to denoise at arbitrary time points within a continuous interval (typically between 0 and 1). This flexibility allows the time discretization strategy for sampling to be selected post-training, enabling the use of advanced sampling techniques (Jolicoeur-Martineau et al., 2021; Zhang & Chen, 2023; Salimans & Ho, 2022; Chung et al., 2022; Song et al., 2020b; Dockhorn et al., 2022) to improve generation performance. These continuous-time frameworks have been successfully extended to graph generative models (Xu et al., 2024; Siraudin et al., 2024), achieving notable improvements.
DeFoG follows a continuous-time formulation, leveraging its flexibility in sampling to achieve enhanced performance while maintaining the strengths of discrete state-space modeling.
A.2.3 From Continuous-Time Discrete Diffusion to Discrete Flow Matching
While both continuous-time discrete diffusion and discrete flow matching (DFM) share the CTMC formulation for the denoising process, they differ fundamentally in the formulation of the noising process. Continuous-time discrete diffusion-based graph generative models (Xu et al., 2024; Siraudin et al., 2024) define the noising process as a CTMC, akin to the denoising process. However, this approach imposes two significant limitations:
  1. 1.
    Incomplete Coupling of Training and Sampling: The rate matrices of the noising and denoising processes are explicitly interrelated, and the noising rate matrix must be fixed during training. This restricts the sampling stage, preventing full decoupling of training and sampling.
  2. 2.
    Limited Design Space: The noising process must be derived analytically, which is not straightforward and is only feasible for rate matrices suitable for matrix exponentiation. Additionally, the denoising rate matrix is implicitly defined during training, constraining the flexibility of the denoising trajectory at sampling time (e.g., fixing the level of stochasticity).
In contrast, DeFoG allows for direct prescription of the noising process, $p_{t|1}$, without these constraints. The rate matrix for the denoising process is selected exclusively at sampling time, fully decoupling the training and sampling stages. This flexibility enables performance optimization during sampling, such as tuning the stochasticity of the denoising trajectory via $R^{\mathrm{DB}}_{t}$ or adjusting target guidance magnitude with $R^{\omega}_{t}$.
The benefits of this decoupled framework are evident in Figures 4, 10 and 12, which demonstrate that the vanilla DeFoG configuration alone does not outperform existing diffusion-based graph generative models. However, our extensive sampling optimization pipeline capitalizes on DeFoG’s flexible design space to achieve state-of-the-art results. These observations align with findings in iterative refinement methods across other data modalities. For instance, Karras et al. (2022) elaborate on the benefits of stochasticity adjustment in denoising trajectories within diffusion models for image generation.
For a comprehensive discussion of the differences between continuous-time discrete diffusion and DFM frameworks, see Appendix H of Campbell et al. (2024).
A.2.4 Mixed Integration of Continuous and Discrete State-spaces
Integrating continuous and categorical data within graph generative models is an important challenge, as many real-world applications involve heterogeneous data types (e.g., molecular graphs containing atomic coordinates alongside categorical atom and bond types). A recent example addressing this challenge is GBD (Liu et al., 2024), which incorporates beta diffusion to jointly model both continuous and discrete variables. Similarly, DeFoG is amenable to formulations involving mixed data types by leveraging an approach akin to MiDi (Vignac et al., 2023), independently factorizing continuous and discrete variables. However, explicitly exploring this integration is beyond the scope of this work.
Appendix B Sample Optimization
@@
In this section, we explore the proposed sampling optimization in more detail. We start by analysing the different time distortion functions in Sec. B.1. Next, in Sec. B.2, we prove that the proposed target guidance mechanism actually satisfies the Kolmogorov equation, thus yielding valid rate matrices and, in Sec. B.3, we provide more details about the detailed balance equation and how it widens the design space of rate matrices. In Sec. B.4, we also describe the adopted sampling optimization pipeline. Finally, in Sec. B.5, we provide more details to better clarify the dynamics of the sampling process.
B.1 Time Distortion Functions
@@
Note that even though we apply the same time distortions for both training and sample stages, in each setting they have different roles: in training, the time distortions skew the PDFs from where $t^{\prime}$ is sampled, while in sampling they vary the denoising step sizes.
B.2 Target Guidance
@@
B.3 Detailed Balance, Prior Incorporation, and Stochasticity
Campbell et al. (2024) show that although their $z_{1}$-conditional formulation of $R^{*}_{t}$ generates $p_{t|1}$, it does not span the full space of valid rate matrices — those that satisfy the conditional Kolmogorov equation (Eq. 8). They derive sufficient conditions for identifying other valid rate matrices. Notably, they demonstrate that matrices of the form